OpenStudy (anonymous):

estamate 4 3/4 - 1 2/3

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

the exact is 37/12

OpenStudy (anonymous):

no... 37/12

OpenStudy (anonymous):

37/12

OpenStudy (anonymous):

see?

OpenStudy (anonymous):

come on 4 3/4 = 4*4+3 = 19 1 2/3 = 3*1 +2 = 5 19/4 - 5/3 19(3)/4(3)-5(4)/3(4) 57/12-20/12=37/12